OpenStudy (anonymous):

100-(x-y)^2 factoring polynomials

OpenStudy (anonymous):

(10-x+y)(10+x-y)

myininaya (myininaya):

\[10^2-(x-y)^2=(10-(x-y))(10+(x-y))\]

myininaya (myininaya):

\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

Thank You